an automobile dealer provides scooter and motorcycles with three body patterns and four different colours find the no of choises open to customer

Grade:12

3 Answers

karthik satyanarayana
4 years ago
as he supplies scoters and motorcycles .we have 12 chooses for scooters and 12 options for motor cycles.
Arun
4 years ago
Dear student
 
total number of choices are = 3 ( for body patterns ) * 4 ( for different colours)
 = 12
 
Hope it helps
 
Incase of any difficulty, please feel free to ask again
Vikas TU
4 years ago
Dear student 
This is AND operation 
No. of vehicle options (Motorcycle, scooter)=2
No. of body patterns=3
No. of different colors=4
No. of options for customer= 4*3*2 = 24